### v1

# Category-free Out-of-Distribution Node Detection with Feature Resonance

Abstract
Detecting out-of-distribution (OOD) nodes in the graph-based machine-learning field is challenging, particularly when in-distribution (ID) node multi-category labels are unavailable. Thus, we focus on feature space rather than label space and find that, ideally, during the optimization of known ID samples, unknown ID samples undergo more significant representation changes than OOD samples, even if the model is trained to fit random targets, which we called the Feature Resonance phenomenon. The rationale behind it is that even without gold labels, the local manifold may still exhibit smooth resonance. Based on this, we further develop a novel graph OOD framework, dubbed Resonance-based Separation and Learning (RSL), which comprises two core modules: (i)-a more practical micro-level proxy of feature resonance that measures the movement of feature vectors in one training step. (ii)-integrate with synthetic OOD nodes strategy to train an effective OOD classifier. Theoretically, we derive an error bound showing the superior separability of OOD nodes during the resonance period. Empirically, RSL achieves state-of-the-art performance, reducing the FPR95 metric by an average of 18.51% across five real-world datasets.
Machine Learning, ICML
1 Introduction
(a) Toy Dataset
(b) Gradient Descent Trajectory
Figure 1: (a) We conduct a preliminary study on the changes in ID and OOD node representations during training using a toy dataset. (b) Projections of the representations of ID and OOD nodes onto gradients: $\text{Proj}_{\nabla\ell(\theta_{t};\cdot)}\mathbf{x}_{i}=\frac{\mathbf{x}_{i}\cdot\nabla\ell(\theta_{t};\cdot)}{\parallel\nabla\ell(\theta_{t};\cdot)\parallel_{2}^{2}}\cdot\nabla\ell(\theta_{t};\cdot)$.
Graph-based machine learning models like Graph Neural Networks (GNNs) (Kipf & Welling, 2016a; Xu et al., 2018; Abu-El-Haija et al., 2019) have become increasingly prevalent in applications such as social network analysis (Fan et al., 2019), knowledge graphs (Baek et al., 2020), and biological networks (De Cao & Kipf, 2018). Despite the success of GNNs, detecting out-of-distribution (OOD) nodes remains an under-explored challenge. These OOD nodes differ significantly from the in-distribution (ID) nodes used during training, and their presence can severely undermine the performance and robustness of graph models. As deploying GNNs in real-world environments becomes more common, the ability to identify and handle OOD nodes is crucial for ensuring the reliability of using these models.
To address this, most existing methods (Hendrycks & Gimpel, 2016; Liang et al., 2017; Hendrycks et al., 2018; Liu et al., 2020; Wu et al., 2023) assume that ID nodes are equipped with multi-category labels. Then, they train an in-distribution classifier and develop OOD metrics based on (i)-classifier outputs, such as Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016) and Energy (Liu et al., 2020; Wu et al., 2023); (ii)-supervised representations, such as KNN (Sun et al., 2022) and NNGuide (Park et al., 2023). However, in many real-world scenarios, accessing multi-category labels for ID nodes is often highly challenging due to practical limitations such as the high cost of annotation, lack of domain expertise, or data privacy concerns, which essentially hinder the effectiveness of current OOD methods. To date, only a few papers (Gong & Sun, 2024; Sun et al., 2022; Sehwag et al., 2021; Liu et al., 2023) study this practical setup, and there is still a large room for improvement, especially in the graph field at the node level.
In this paper, we revisit the graph OOD task at the node level from a new perspective and turn our attention to the intrinsic similarities within the data. An intuitive idea is that the ID samples may still share some commonalities in the representation space. We hypothesize that when optimizing the representation of known ID nodes, the representation of unknown ID nodes and unknown OOD nodes will change with different trajectories. Based on the hypothesis and using a toy dataset (Figure 1(a)), we design an experiment where the features of labeled ID samples are aligned to an arbitrarily fixed representation vector. Interestingly, we observe a distinct behavior during this optimization process: the representations of unlabeled wild ID samples experienced more pronounced changes than wild OOD samples, as shown in Figure 1(b). This phenomenon closely resembles the concept of forced vibration, where resonance occurs when an external force aligns with the natural frequency of an oscillator, amplifying its oscillation to a maximum. Analogously, we refer to this phenomenon as Feature Resonance: during the optimization of known ID samples, the representation of unknown ID samples undergoes more significant changes compared to OOD samples. This phenomenon reveals the intrinsic relationship between ID samples, highlighting their shared underlying distribution. Evidently, this feature resonance phenomenon can be leveraged for OOD detection: weaker representation changes during known ID optimization indicate a higher likelihood of being OOD.
In real-world scenarios, due to the intrinsic complex pattern in data, we find that the feature resonance phenomenon still occurs but slightly differs from the ideal conditions. To illustrate this, we further propose a micro-level proxy for measuring feature resonance—by computing the movement of the representation vector in one training step. Our findings reveal that in more complex scenarios, the feature resonance phenomenon typically arises during the middle stages of the training process, whereas during other phases, it may be overwhelmed by noise or obscured by overfitting. In such cases, evaluating the entire trajectory often fails to yield satisfactory results. Fortunately, efficient OOD detection can still be achieved by calculating the micro-level feature resonance measure. By utilizing a simple binary ID/OOD validation set [^1] The use of the validation set is consistent with previous works (Katz-Samuels et al., 2022; Gong & Sun, 2024; Du et al., 2024a, b) and does not contain multi-category labels., we empirically show the feature resonance period can be precisely identified, and we identify more minor representation differences as OOD samples. Notably, our new micro-level feature resonance measure is still label-independent by fitting a randomly fixed target, making it highly compelling in category-free scenarios. Theoretical and experimental proof that micro-level feature resonance can filter a set of reliable OOD nodes with low error, e.g., on the FPR95 metric, the micro-level feature resonance achieves an average reduction of 10.93% compared to current state-of-the-art methods.
Furthermore, we combine the micro-level feature resonance with the current Langevin-based synthetic OOD nodes generating strategy to train an OOD classifier for more effective OOD node detection performance, which we call the whole framework as RSL; for example, the FPR95 metric is reduced by an average of 18.51% compared to the current state-of-the-art methods.
2 Preliminaries
Graph Neural Network. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ denote an graph, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of nodes and edges, respectively. We represent the node space by $\mathbb{V}$, and $\mathcal{V}\in\mathbb{V}$. ${\boldsymbol{X}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ denote the matrix of features of the nodes. Here, the representation of a node $v$ can be defined as $\mathbf{h}_{v}$. Graph neural networks (GNNs) aim to update the representation of the given graph $\mathcal{G}$ by leveraging its topological structure. For the representation $\mathbf{h}_{v}$ of node $v$, its propagation of the $k$-th layer GNN is represented as:
$\begin{array}[]{cl}\mathbf{h}_{v}^{(k)}&=g(\mathbf{h}_{v}^{(k-1)};\theta)\\ &=\mathrm{UP}^{(k)}\big{(}\mathrm{AGG}^{(k)}\big{(}\mathbf{h}_{u}^{(k-1)}:\forall u\in\mathcal{N}(v)\cup v\big{)}\big{)}\end{array}$ (1)
where $g(\cdot;\theta)$ denotes the GNN encoder, $\theta$ represents all trainable parameters of GNN encoder. $\mathrm{AGG}(\cdot)$ denotes a function that aggregates messages from the neighbors of node $v$, $\mathcal{N}(v)$ represents the set of neighbors. $\mathrm{UP}(\cdot)$ denotes a function that updates the representation of node $v$ with the current representation of $v$ and the aggregated vector.
Problem Statement. The node set can be divided into $\mathcal{V}=\mathcal{V}_{\mathrm{in}}\cup\mathcal{V}_{\mathrm{out}}$, where $\mathcal{V}_{\mathrm{in}}$ and $\mathcal{V}_{\mathrm{out}}$ represents the ID node set and OOD node set, respectively. We assume ID nodes are sampled from the distribution $\mathbb{P}_{\mathrm{in}}$, and OOD nodes are sampled from distribution $\mathbb{P}_{\mathrm{out}}$. We formally define the Category-free OOD node detection task:
Definition 2.1.
Category-free OOD node detection. Given a collection of nodes sampled from $\mathbb{P}_{\mathrm{in}}$ and $\mathbb{P}_{\mathrm{out}}$, the objective is to correctly identify the source of each node, whether it is from the $\mathbb{P}_{\mathrm{in}}$ or $\mathbb{P}_{\mathrm{out}}$.
Unlabeled Wild Node. In this work, we incorporate unlabeled wild node $\mathcal{V}_{\mathrm{wild}}=\{\tilde{v}_{1},\cdots,\tilde{v}_{m}\}$ with feature $\boldsymbol{X}_{\mathrm{wild}}=\{\tilde{\mathbf{x}}_{1},\cdots,\tilde{\mathbf{x}}_{m}\}$ into our learning framework, leveraging the fact that such features are often accessible. We define the unlabeled wild nodes distribution as follows:
Definition 2.2.
Unlabeled wild nodes. Unlabeled wild nodes typically consist of a mixture of ID nodes and OOD nodes. We use the Huber contamination model (Huber, 1992) to characterize the marginal distribution of the wild data:
$\mathbb{P}_{\mathrm{wild}}=(1-\pi)\mathbb{P}_{\mathrm{in}}+\pi\mathbb{P}_{\mathrm{out}}$ (2)
where $\pi\in(0,1]$.
Figure 2: Schematic of Feature Resonance.
3 Method
3.1 Revealing the Feature Resonance Phenomenon
Previous studies (Hendrycks & Gimpel, 2016; Liu et al., 2020; Wu et al., 2023) mostly train a classifier on ID nodes with multi-category labels and develop selection criteria based on output probabilities, e.g. entropy. However, these methods become inapplicable in category-free scenarios.
To address this problem, we turn our attention to the intrinsic similarities within the data. An intuitive idea is that although the output space may no longer be reliable, the ID samples may still share some commonalities in the representation space. We hypothesize that when optimizing the representation of known ID nodes, the representation of unknown ID nodes and unknown OOD nodes will change with different trajectories. Motivated by this, and under the assumption of some specific training process, we define a feature trajectory measure $\hat{F}(\tilde{\mathbf{x}}_{i})$ of a sample $\tilde{\mathbf{x}}_{i}$:
$\hat{F}(\tilde{\mathbf{x}}_{i})=\sum_{t}h_{\theta_{t+1}}(\tilde{\mathbf{x}}_{i})-h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})$ (3)
where $h_{\theta_{t}}$ denote the model that performs representation transformation on a sample $\tilde{\mathbf{x}}_{i}$, with $\theta_{t}$ representing its parameters at the $t$-th epoch.
[…]
Despite the promise, the abovementioned feature resonance phenomenon occurs under multi-category training. But how can we induce this phenomenon in a category-free scenario without multi-category labels? Interestingly, we find that even when random labels are assigned to known ID nodes for multi-category training, the trajectories of unknown ID nodes are still more significant than those of unknown OOD nodes. More surprisingly, on a ideal toy dataset, even when all known ID node representations are aligned toward one single random fixed target vector, the trajectories of unknown ID nodes are still longer than those of unknown OOD nodes, as shown in Figure 1. Green points represent unknown ID samples, blue points represent unknown OOD samples, and red points represent known ID samples aligned to a target vector. As shown in Figure 1(b), modifying the representation of known ID samples results in longer representation change trajectories for unknown ID samples compared to unknown OOD samples. The experiments above indicate that the feature resonance phenomenon is label-independent and results from the intrinsic relationships between ID node representations. Therefore, this is highly suitable for category-free OOD detection scenarios without multi-category labels.
Since the trajectory represents a global change, we call it a macroscopic feature resonance, as follows:
Definition 3.1.
Feature Resonance (macroscopic): For any optimization objective $\ell(\boldsymbol{X}_{\text{known}},\cdot)$ applied to the representations $\boldsymbol{X}_{\text{known}}$ of known ID samples derived from any model $h_{\theta}(\cdot)$, we have $\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$.
3.2 Utilizing the Micro-level Feature Resonance Phenomenon with An Arbitrary Target
As mentioned above, we can leverage the feature resonance phenomenon to detect OOD nodes. In our realistic implementations, we align the features of known ID nodes to an arbitrary target vector using mean squared error as follows:
$\ell(h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{1}^{\top}e-({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (4)
where $h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}})={\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top}$ represent the last linear layer of the model for representation transformation and $e$ denotes an arbitrary randomly generated target vector.
But, in contrast to our toy dataset, the real-world datasets typically exhibit much more complex feature attributes. As a result, the feature resonance of trajectory at the macro level is not as ideal or pronounced as observed in experiments on the toy dataset. Therefore, to explore the reasons behind this issue, we delve deeper into the changes in finer-grained node representations across epochs to study the feature resonance phenomenon. Specifically, we study the differences in $\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})=h_{\theta_{t+1}}(\tilde{\mathbf{x}}_{i})-h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})$ between ID samples and OOD samples. Obviously, the existence of $\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$ is a necessary condition for satisfying $\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$, so we define $\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$ as a feature resonance at the microscopic level:
Definition 3.2.
Feature Resonance (microscopic): For any optimization objective $\ell(\boldsymbol{X}_{\text{known}},\cdot)$ applied to the known ID nodes’ representations $\boldsymbol{X}_{\text{known}}$ from any model $h_{\theta_{t}}(\cdot)$, during the optimization process, there exists $t$ such that $\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$. We define the resonance-based filtering score as $\tau_{i}=\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{2}$. The resonance-based scores $\tau$ of OOD nodes should be smaller than those of ID nodes at $t$.
By observing $\tau$ for ID samples and OOD samples, we find that feature resonance does not persist throughout the entire training process but rather occurs at specific stages of training. In our experiments on the common benchmarks, we find that during the early stages of training, the model is searching for the optimal optimization path, leading to chaotic representation changes and thus making feature resonance insignificant. However, in the middle stages of training, once the model identifies an optimization path that aligns with the patterns of the ID samples, it optimizes along the path most relevant to the features of the ID samples, and feature resonance becomes most prominent. As the model continues to optimize and enters the overfitting stage, the feature resonance phenomenon begins to dissipate. Figure 3 shows the experimental results on the Amazon dataset, and others are provided in Figure 6 of the Appendix. Through the above experiments and analyses, we find that using $\hat{F}(\tilde{\mathbf{x}}_{i})$ to identify OOD nodes is affected by error accumulation and is, therefore, not a reliable approach. However, there exists a specific period during training when micro-level feature resonance occurs. By utilizing a validation set (Katz-Samuels et al., 2022; Gong & Sun, 2024; Du et al., 2024a, b), we can easily identify the period during which feature resonance occurs.
Formally, our new feature resonance-based OOD nodes detector is defined as follows:
$\begin{split}&g_{\gamma}(\tilde{\mathbf{x}}_{i})=\mathds{1}\{\tau_{i}^{*}\leq\gamma\},\\ &\text{s.t.},\tau^{*}=\max_{t}\text{AUROC}(\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{in}}},\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{out}}})\end{split}$ (5)
where $g_{\gamma}=1$ indicates the OOD nodes while $g_{\gamma}=0$ indicates otherwise, and $\gamma$ is typically chosen to guarantee a high percentage, such as 95%, of ID data that is correctly classified. Here, $t$ is determined by the validation set $\mathcal{V}_{\text{val}}$.
Figure 3: The performance of using resonance-based score $\tau$ to detect OOD nodes varies with training progress. The higher the AUROC, the better, and the lower the FPR95, the better.
To summarize our method: we calculate a resonance-based filtering score $\tau$ during the transformation of known ID sample representations. By leveraging a validation set, we identify the period during training when micro-level resonance is most significant. Within this period, test set nodes with smaller $\tau$ values are more likely to be OOD nodes.
3.3 Extension with Synthetic OOD Node Strategy
Although the resonance-based filtering score effectively separates OOD nodes, recent studies (Gong & Sun, 2024) suggest that training an OOD classifier with synthetic OOD nodes can improve OOD node detection. Therefore, we propose a novel framework that employs feature resonance scores to generate more realistic synthetic OOD nodes.
Specifically, we define the candidate OOD node set as $\mathcal{V}_{\mathrm{cand}}=\{\tilde{v}_{i}\in\mathcal{V}_{\mathrm{wild}}:\tau_{i}\leq T\}$, where $T=\mathrm{min}_{\mathrm{n}}(\tau)$ is the $n$-th smallest $\tau$ of wild nodes, selecting nodes with the smallest $n$ $\tau$ values. The features of these nodes form $\boldsymbol{X}_{\mathrm{cand}}$. Then, we compute a trainable metric based on the weighted mapping of node $v$’s representations across $K$ GNN layers: $E_{\theta}(v)=\mathbf{W}_{K}\big{(}\sum_{k}^{K}\beta_{k}\mathbf{h}_{v}^{(k)}\big{)}$, where $\beta_{k}\in\mathbb{R}$ is a learnable parameter, and $\mathbf{W}_{K}\in\mathbb{R}^{1\times d}$ transforms the node representations to the energy scalar. Then, we employ stochastic gradient Langevin dynamics (SGLD) (Welling & Teh, 2011) to generate synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}=\{\hat{v}_{1},\cdots,\hat{v}_{j}\}$ with random initial features $\boldsymbol{X}_{\mathrm{syn}}=\{\hat{\mathbf{x}}_{1},\cdots,\hat{\mathbf{x}}_{j}\}$ as follows:
$\begin{split}\hat{\mathbf{x}}_{j}^{(t+1)}=&\lambda\big{(}\hat{\mathbf{x}}_{j}^{(t)}-\frac{\alpha}{2}\nabla_{\hat{\mathbf{x}}_{j}^{(t)}}E_{\theta}\big{(}\hat{v}_{j}^{(t)}\big{)}+\epsilon\big{)}\\ &+(1-\lambda)\mathbb{E}_{\mathbf{x}\sim\boldsymbol{X}_{\mathrm{cand}}}(\mathbf{x}-\hat{\mathbf{x}}_{j}^{(t)})\end{split}$ (6)
where $\frac{\alpha}{2}$ is the step size and $\lambda$ is a trade-off hyperparameter. $\epsilon$ is the Gaussian noise sampled from multivariate Gaussian distribution $\mathcal{N}(0,\zeta)$. Unlike EnergyDef (Gong & Sun, 2024), we utilize the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ as examples to generate synthetic OOD nodes that better align with the actual OOD nodes. After obtaining the synthetic OOD nodes, we define the training set $\mathcal{V}_{\mathrm{train}}=\mathcal{V}_{\text{known}}\ \cup\ \mathcal{V}_{\mathrm{cand}}\ \cup\ \mathcal{V}_{\mathrm{syn}}$ with features $\boldsymbol{X}_{\mathrm{train}}$ and labels $\boldsymbol{Y}_{\mathrm{train}}$. The initially known ID nodes $\mathcal{V}_{\mathrm{known}}$ are assigned a label of $1$. In contrast, the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ and the generated synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}$ are assigned a label of 00. We use binary cross-entropy loss for training:
$\ell_{\text{cls}}=-\big{(}{\mathrm{y}}_{v}\mathrm{log}(\sigma(E_{\theta}(v)))+{(1-\mathrm{y}}_{v})\mathrm{log}(1-\sigma(E_{\theta}(v)))\big{)}$ (7)
where $\sigma(\cdot)$ is the sigmod function. Similarly, we identify the OOD nodes as follows: $g^{\prime}_{\gamma^{\prime}}(E_{\theta}(v))=\mathds{1}\{E_{\theta}(v)\leq\gamma^{\prime}\}$. , where $g^{\prime}_{\gamma^{\prime}}=1$ indicates the OOD nodes while $g^{\prime}_{\gamma^{\prime}}=0$ indicates otherwise, and $\gamma^{\prime}$ is chosen to guarantee a high percentage, e.g., 95%, of ID data that is correctly classified.
Table 1: The statistics of the real-world OOD node detection datasets. $\times$ denotes no available multi-category labels. Notably, even on Squirrel and WikiCS, we do not use any true label as well.
Dataset Squirrel WikiCS YelpChi Amazon Reddit
# Nodes 5,201 11,701 45,954 11,944 10,984
# Features 2,089 300 32 25 64
Avg. Degree 41.7 36.9 175.2 800.2 15.3
OOD node (%) 20.0 29.5 14.5 9.5 3.3
# Category 5 10 $\times$ $\times$ $\times$
Table 2: Category-free OOD detection on real-world datasets. “OOM” indicates out-of-memory, “TLE” means time limit exceeded, and “-” denotes inapplicability. Detectors with $\clubsuit$ use only node attributes, while $\spadesuit$ share RSL’s GNN backbone. Entropy-based methods with $\lozenge$ use true multi-category labels, and $\blacklozenge$ rely on K-means pseudo labels. Top results: 1st, 2nd.
Method Dataset Squirrel WikiCS YelpChi Amazon Reddit
[…]
$\text{MLPAE}^{\clubsuit}$ 43.15 24.81 97.98 70.99 63.74 77.76 51.90 24.53 92.42 74.54 51.59 57.93 52.10 5.80 94.43
GCNAE 37.87 22.64 99.08 57.95 46.32 92.97 44.20 19.22 97.06 45.07 12.38 98.54 51.78 6.14 93.75
GAAN 38.01 22.57 98.99 58.15 46.60 93.37 44.29 19.30 96.91 53.26 6.63 98.05 52.21 5.96 94.06
DOMINANT 41.78 24.73 95.53 42.55 35.43 97.22 52.77 24.90 92.86 78.08 35.96 76.05 55.89 6.03 96.48
ANOMALOUS 51.04 29.09 96.39 67.99 54.51 92.74 OOM OOM OOM 65.12 25.15 85.34 55.18 6.40 94.10
SL-GAD 48.29 27.62 97.19 51.87 44.83 95.26 56.11 26.49 93.27 82.63 56.27 51.36 51.63 6.02 94.27
$\text{GOAD}^{\spadesuit}$ 62.32 37.51 92.28 50.65 37.22 99.78 58.03 28.51 89.84 72.92 45.53 66.36 52.89 5.36 94.26
$\text{NeuTral AD}^{\spadesuit}$ 52.51 30.04 97.16 53.58 43.49 94.30 55.81 25.14 94.23 70.01 24.36 92.19 55.70 6.45 94.59
$\text{GKDE}^{\lozenge}$ 56.15 33.41 94.96 70.47 61.18 82.71 - - - - - - - - -
$\text{OODGAT}^{\lozenge}$ 58.84 35.13 93.31 74.13 62.47 84.48 - - - - - - - - -
$\text{GNNSafe}^{\spadesuit\lozenge}$ 56.38 32.22 95.17 73.35 66.47 76.24 - - - - - - - - -
$\text{OODGAT}^{\blacklozenge}$ 57.78 34.66 92.61 52.76 44.71 90.02 55.97 23.07 97.93 82.54 54.94 52.10 54.62 6.05 93.85
$\text{GNNSafe}^{\spadesuit\blacklozenge}$ 49.52 26.63 97.60 64.15 50.85 92.63 55.26 26.68 91.40 68.51 25.39 84.31 49.63 5.36 95.98
$\text{SSD}^{\spadesuit}$ TLE TLE TLE 64.29 58.45 87.12 55.39 27.88 91.63 72.49 41.82 84.27 59.74 6.21 91.15
$\text{Energy{Def}}^{\spadesuit}$ 64.15 37.40 91.77 70.22 60.10 83.17 62.04 29.71 90.62 86.57 74.50 32.43 63.32 8.34 89.34
RSL w/o classifier 61.52 38.96 90.18 79.15 78.65 70.38 65.42 37.08 83.53 87.43 83.31 19.56 52.37 6.97 91.39
RSL w/o $\mathcal{V}_{\mathrm{syn}}$ 60.46 34.89 93.59 81.21 79.93 52.19 65.15 38.93 81.84 87.81 81.10 25.18 61.36 8.48 89.43
RSL 64.12 39.58 89.90 84.01 81.14 49.23 66.11 39.73 80.45 90.03 83.91 19.60 64.83 10.18 85.49
Table 3: The effectiveness of different OOD candidate node selection strategies.
Method Dataset Squirrel WikiCS YelpChi Amazon Reddit
[…]
RSL w/ Cosine Similarity 64.00 38.11 91.46 81.61 76.36 70.38 59.76 35.03 85.89 83.35 74.85 27.63 54.07 7.25 92.21
RSL w/ Euclidean Distance 64.01 39.30 90.45 78.63 74.28 63.26 52.53 24.20 93.53 53.08 18.29 93.64 62.19 8.38 90.90
RSL w/ Mahalanobis Distance TLE TLE TLE 83.18 79.11 58.03 54.07 25.44 92.40 63.71 30.66 79.96 60.81 8.42 90.08
RSL w/ EnergyDef 63.66 38.29 91.69 61.21 50.41 90.42 57.33 26.79 91.90 77.72 55.23 54.52 61.90 8.55 89.51
RSL w/ Resonance-based Score $\tau$ 64.12 39.58 89.90 84.01 81.14 49.23 66.11 39.73 80.45 90.03 83.91 19.60 64.83 10.18 85.49
3.4 Theoretical Analysis
Our main theorem quantifies the separability of the outliers in the wild by using the resonance-based filter score $\tau$. We provide detailed theoretical proof in the Appendix C.
Let $\text{ERR}_{\text{out}}^{t}$ be the error rate of OOD data being regarded as ID at $t$-th epoch, i.e., $\text{ERR}_{\text{out}}^{t}=|\{\tilde{v}_{i}\in\mathcal{V}_{\text{wild}}^{\text{out}}:\tau_{i}\geq T\}|/|\mathcal{V}_{\text{wild}}^{\text{out}}|$, where $\mathcal{V}_{\text{wild}}^{\text{out}}$ denotes the set of outliers from the wild data $\mathcal{V}_{\text{wild}}$. Then $\text{ERR}_{\text{out}}$ has the following generalization bound:
Theorem 3.3.
(Informal). Under mild conditions, if $\ell(\mathbf{x},e)$ is $\beta$-smooth w.r.t $\mathbf{w}_{t}$, $\mathbb{P}_{\mathrm{wild}}$ has $(\gamma,\xi)$-discrepancy w.r.t $\mathbb{P}_{\mathrm{in}}$, and there is $\eta\in(0,1)$ s.t. $\Delta=(1-\eta)^{2}\xi^{2}-8\beta_{1}R_{in}^{*}>0$, then where $n=\Omega(d/\mathrm{min}\{\eta^{2}\Delta,(\gamma-R_{in}^{*})\}),m=\Omega(d/\eta^{2}\xi^{2})$, with the probability at least 0.9, for $0<T<0.9\widehat{M}_{t}$($\widehat{M}_{t}$ is the upper bound of score $\tau_{i}$),
$\begin{split}\text{ERR}_{\text{out}}^{t}\leq&\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\\ &+O(\sqrt{\frac{d}{\pi^{2}n}})+O(\sqrt{\frac{\mathrm{max}\{d,\Delta_{\xi}^{\eta^{2}}/\pi^{2}\}}{\pi^{2}(1-\pi)m}})\end{split}$ (8)
where $\Delta_{\xi}^{\eta}=0.98\eta^{2}\xi^{2}-8\beta_{1}R_{in}^{*}$ and $R_{in}^{*}$ is the optimal ID risk, i.e., $R_{in}^{*}=\mathrm{min}_{\mathbf{w}\in\mathcal{W}}\mathbb{E_{\mathbf{x}\sim\mathbb{P}_{\mathrm{in}}}}\ell(\mathbf{x},e)$. $d$ is the dimension of the space $\mathcal{W}$, $t$ denotes the $t$-th epoch, and $\pi$ is the OOD class-prior probability in the wild.
Practical implications of Therorem 3.3. The above theorem states that under mild assumptions, the error $ERR_{\mathrm{out}}$ is upper bounded. If the following two regulatory conditions hold: 1) the sizes of the labeled ID $n$ and wild data $m$ are sufficiently large; 2) the optimal ID risk $R_{in}^{*}$ is small, then the upper bound is mainly depended on $T$ and $t$. We further study the main error of $T$ and $t$ which we defined as $\delta(T,t)$.
Theorem 3.4.
(Informal). 1) if $\Delta_{\xi}^{\eta}\geq(1-\epsilon)\pi$ for a small error $\epsilon\geq 0$, then the main error $\delta(T,t)$ satisfies that
$\begin{split}\delta(T,t)&=\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\\ &\leq\frac{\epsilon}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\end{split}$ (9)
2) When learning rate $\alpha$ is small sufficiently, and if $\xi\geq 2.011\sqrt{8\beta_{1}R_{in}^{*}+1.011\sqrt{\pi}}$, then there exists $\eta\in(0,1)$ ensuring that $\Delta>0$ and $\Delta_{\xi}^{\eta}>\pi$ hold, which implies that the main error $\delta(T,t)=0$.
Practical implications of Therorem 3.4. Theorem 3.4 states that when the learning rate $\alpha$ is sufficiently small, the primary error $\delta(T,t)$ can approach zero if the difference $\zeta$ between the two data distributions $\mathbb{P}_{\text{wild}}$ and $\mathbb{P}_{\text{in}}$ is greater than a certain small value. Meanwhile, Theorem 3.4 also shows that the primary error $\delta(T,t)$ is inversely proportional to the learning rate $\alpha$ and the number of epochs ($t$). As the $t$ increases, the primary error $\delta(T,t)$ also increases, while a smaller learning rate $\alpha$ leads to a minor primary error $\delta(T,t)$. However, during training, there exists $t$ at which the error reaches its minimum.
4 Experiment
4.1 Experimental Setup
Datasets.
We conduct extensive experiments to evaluate RSL on five real-world OOD node detection datasets: Two multi-category datasets, Squirrel (Rozemberczki et al., 2021) and WikiCS (Mernyei & Cangea, 2020), and three binary classification fraud detection datasets: YelpChi (Rayana & Akoglu, 2015), Amazon (McAuley & Leskovec, 2013), and Reddit (Kumar et al., 2019). The statistics of these datasets are summarized in Table 1. We provide detailed dataset description in the Appendix F.3.
Figure 4: Performance of detecting OOD nodes with different metrics. $\tau$ represents the resonance-based score, the “Overall Trajectory” represents the total cumulative length of the training trajectory $\hat{F}(\tilde{\mathbf{x}}_{i})=\sum_{t}\tau_{i}$, and the “Sliding Window” refers to the cumulative $\tau$ within a window of width 10: $\hat{F}_{10}(\tilde{\mathbf{x}}_{i})=\sum^{t}_{t-10}\tau_{i}$.
Baselines.
We assess the performance of RSL against a diverse range of baseline methods spanning five categories: 1) Traditional outlier detection methods, including local outlier factor (Breunig et al., 2000) with $k$-nearest neighbors (LOF-KNN) and MLP autoencoder (MLPAE). 2) Graph-based outlier detection models, including GCN autoencoder (Kipf & Welling, 2016b), GAAN (Chen et al., 2020), DOMINANT (Ding et al., 2019), ANOMALOUS (Peng et al., 2018), and SL-GAD (Zheng et al., 2021). 3) Transformation-based outlier detection approaches, such as GOAD (Bergman & Hoshen, 2020) and NeuTral AD (Qiu et al., 2021). 4) Entropy-based detection techniques, including GKDE (Zhao et al., 2020), OODGAT (Song & Wang, 2022), and GNNSafe (Wu et al., 2023). 5) Category-free detection methods, including EnergyDef (Gong & Sun, 2024) and SSD (Sehwag et al., 2021). Details of baselines and implementation are in Appendix F.4 and F.5, respectively.
(a) EnergyDef
(b) Resonance-based Filter Score $\tau$ (Ours)
(c) RSL (Ours)
Figure 5: The score distribution of ID nodes and OOD nodes on Amazon obtained using different methods.
Metrics.
Following prior research on OOD node detection, we evaluate the detection performance using three widely recognized, threshold-independent metrics: AUROC ($\uparrow$), AUPR ($\uparrow$) and FPR95($\downarrow$). We provide a detailed metric description in the Appendix F.2.
4.2 Main Results
Table 2 presents the main experimental results of various methods across five public datasets. The traditional methods like LOF-KNN and MLPAE perform poorly across most datasets, particularly with high false positive rates. Graph-based methods such as GCNAE, GAAN, and DOMINANT show some decent results but generally lag behind RSL and EnergyDef. Entropy-based methods like OODGAT and GNNSafe perform well on datasets with multi-category label information (e.g., WikiCS) but struggle on datasets without such labels, like YelpChi. Overall, these methods tend to be less robust compared to RSL. Specifically, RSL achieves significant improvements on most datasets. On average, it improves by 5.24% and 19.26% in AUROC and AUPR, respectively, and reduces FPR95 by 18.51%. Notably, when additional multi-category label information is available, GNNSafe outperforms EnergyDef on WikiCS. However, even in the absence of multi-class label information, our method still outperforms GNNSafe, achieving a remarkable improvement of 13.33% in AUROC, 22.07% in AUPR, and a 35.43% reduction in FPR95. These results powerfully demonstrate the effectiveness of our method.
Table 4: The effectiveness of the resonance-based filter score $\tau$ in filtering OOD nodes with different alignment targets for known ID node representations. True multi-label means aligning ID node representations with multiple target vectors based on true multi-class labels. Multiple random vectors means aligning ID node representations with random target vectors. A random vector means aligning ID node representations with a single target vector.
Method Dataset Target Squirrel WikiCS
[…]
EnergyDef - 64.15 37.40 91.77 70.22 60.10 83.17
RSL w/o classifier True multi-label 61.63 37.12 90.62 71.03 72.47 81.96
RSL w/o classifier Multiple random vectors 61.44 37.39 90.62 73.64 74.13 69.25
RSL w/o classifier A random vector 61.52 38.96 90.18 79.15 78.65 70.38
4.3 Ablation Study
How effective is resonance-based filter score $\tau$?
The experimental results in the row labeled “RSL w/o classifier” of Table 2 show that using the raw resonance-based score $\tau$ to filter OOD nodes is already more effective than the SOTA method on most datasets. On the FPR95 metric, the resonance-based score achieves an average reduction of 10.93% compared to current SOTA methods. Notably, even when compared to methods that leverage additional multi-category label information, our approach continues to demonstrate a clear performance advantage. For example, on the WikiCS dataset, the resonance-based score reduces the FPR95 metric by 7.69% compared to GNNSafe.
How effective are the synthetic OOD nodes combined with the feature resonance score?
The experimental results in the row labeled “RSL w/o $\mathcal{V}_{\mathrm{syn}}$” of Table 2 show that after removing the synthetic OOD nodes, the performance of the trained OOD classifier declined to varying degrees. This indicates that synthetic OOD nodes enhance the generalization ability of the OOD classifier, allowing it to detect more OOD nodes more accurately. It is worth noting that our synthetic OOD nodes, generated by leveraging real OOD nodes selected using $\tau$, better align with real-world OOD scenarios and, therefore, outperform EnergyDef.
4.4 Effectiveness of Feature Resonance Score in Selecting OOD Nodes
We aim to evaluate the performance of RSL when integrated with methods other than the resonance-based score for selecting reliable OOD nodes. To ensure fairness, we used the same parameters and selected the same number of OOD nodes. From a metric learning perspective, we computed the cosine similarity, Euclidean distance, and Mahalanobis distance between unknown nodes and the prototypes of known ID nodes, with smaller values indicating a higher likelihood of being OOD nodes. We also applied EnergyDef for OOD node selection. The results, presented in Table 3, show that, under the same conditions, the OOD nodes selected using $\tau$ are more reliable than those selected by the other methods.
4.5 Effectiveness of Different Scoring Strategies Based on Feature Resonance
We evaluate the effectiveness of three score design strategies based on feature resonance: the resonance-based score $\tau$, the global trajectory norm, and the sliding window accumulation (width 10). As shown in Figure 4, $\tau$ outperforms the other two scores on most datasets. The sliding window approach performs better than the global trajectory norm, with further details on width in Appendix G.3. This indicates that finer-grained information improves OOD node detection, so we select $\tau$ as the primary score for filtering OOD nodes in our method.
4.6 Feature Resonance with Different Target Vectors
We explore micro-level feature resonance using different target vectors through experiments on Squirrel and WikiCS datasets with true $N$-category labels. Based on neural collapse theory (Papyan et al., 2020; Zhou et al., 2022), we set $N$ target vectors that form a simplex equiangular tight frame  [^3] The definition of the simplex equiangular tight frame is introduced in Appendix G.1., maximizing separation. As shown in Table 4, the ”True multi-label” row demonstrates the effectiveness of this approach. Interestingly, even when random labels are assigned (the ”Multiple random vectors” row) or when all ID representations align with a fixed vector (the ”A random vector” row), unknown ID nodes still show larger $\tau$ than unknown OOD nodes, as seen in Table 4. These results suggest that feature resonance is label-independent, stemming from intrinsic relationships between ID node representations, making it suitable for category-free OOD detection.
Table 5: Time cost (s).
Method Dataset Squirrel WikiCS YelpChi Amazon Reddit
EnergyDef 10.94 27.11 76.51 33.81 26.44
RSL w/o classifier 5.25 4.03 5.41 5.75 3.71
RSL 11.54 17.53 74.83 36.33 38.23
4.7 Time Efficiency
We compare the time consumption of our method, RSL, with the current SOTA method, EnergyDef. The experimental results are shown in Table 5. The experiments show that the overall time efficiency of RSL is comparable to that of EnergyDef, with similar time consumption across different datasets. However, it is worth noting that when we use the resonance-based score $\tau$ alone for OOD node detection, its efficiency improves significantly over EnergyDef, with an average reduction of 79.81% in time consumption. This indicates that $\tau$ not only demonstrates significant effectiveness in detecting OOD nodes but also offers high efficiency.
4.8 Score Distribution Visualization
We visualize the score distributions of ID and OOD nodes on the Amazon dataset obtained using different methods, as shown in Figure 5. When using the resonance-based score (Figure 5 (b)), the majority of unknown ID nodes show more significant representation changes compared to unknown OOD nodes. This separation of OOD nodes already exceeds EnergyDef (Figure 5 (a)). After training with synthetic OOD nodes (Figure 5 (c)), the separation between the energy scores of ID and OOD nodes still improves compared to EnergyDef, which demonstrates the effectiveness of RSL.
5 Related Works
General OOD Detection Methods.
OOD detection methods are generally categorized into entropy-based, density-based, and representation-based approaches. Entropy-based methods such as Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016), Energy (Liu et al., 2020), and other methods (Liang et al., 2017; Bendale & Boult, 2016; Hendrycks et al., 2018; Geifman & El-Yaniv, 2019; Malinin & Gales, 2018; Jeong & Kim, 2020; Chen et al., 2021; Wei et al., 2021; Ming et al., 2022b, a) compute OOD scores from class distributions. Still, they rely heavily on multi-category labels, which limits their use in category-free settings. Density-based methods, such as Mahalanobis distance (Lee et al., 2018) and residual flow models (Zisselman & Tamar, 2020), estimate sample probabilities based on their distribution but struggle with handling high-dimensional data and complex relationships (Ren et al., 2019; Serrà et al., 2019). Representation-based methods, including KNN (Sun et al., 2022) and NNGuide (Park et al., 2023), focus on differentiating OOD and ID nodes by analyzing learned embeddings in feature space. However, they still need a pre-trained multi-category ID classifier. In contrast, SSD (Sehwag et al., 2021) is an outlier detector that leverages self-supervised representation learning and Mahalanobis distance-based detection on unlabeled ID data.
Category-free OOD Detection in Graphs.
Category-free OOD detection in graphs aims to identify OOD nodes without relying on multi-category labels, posing unique challenges for traditional methods. Entropy-based methods, such as OODGAT (Song & Wang, 2022) and GNNSafe (Wu et al., 2023), depend on classifier outputs and are not suitable for category-free settings. Representation-based methods, including EnergyDef (Gong & Sun, 2024), aim to generate synthetic OOD nodes but often fail to capture the true features of real OOD nodes. Graph anomaly detection methods, like DOMINANT (Ding et al., 2019) and SL-GAD (Zheng et al., 2021), detect general anomalies through reconstruction errors, but they struggle to distinguish between OOD nodes and general anomalies. Recent works such as (Li et al., 2022; Bazhenov et al., 2022; Liu et al., 2023; Ding & Shi, 2023) explore graph-level OOD detection but can not be directly applied to node-level OOD detection due to the complexity of node dependencies.
6 Conclusion
In this paper, we introduce the concept of Feature Resonance for category-free OOD detection, demonstrating that unknown ID samples undergo more substantial representation changes compared to OOD samples during the optimization of known ID samples, even in the absence of multi-category labels. To effectively capture this phenomenon, we propose a label-independent, micro-level proxy that measures feature vector movements in a single training step. Building on this, we present the RSL framework, which integrates the micro-level feature resonance with synthetic OOD node generation via SGLD, enhancing OOD detection performance and offering an efficient and practical solution for category-free OOD node detection.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Notations, Definitions, Assumptions and Important Constants
A.1 Notations
Notation Description
Spaces
$\boldsymbol{X}$, $\boldsymbol{Y}$ the input space and the label space.
$\mathcal{W}$ the hypothesis spaces.
Distributions
$\mathbb{P}_{\text{wild}},\mathbb{P}_{\text{in}},\mathbb{P}_{\text{out}}$ data distribution for wild data, labeled ID data and OOD data.
$\mathbb{P}_{\boldsymbol{XY}}$ the joint data distribution for ID data..
Data and Models
$\mathbf{w},\mathbf{x}$ weight, input.
^$\widehat{\nabla},\tau$ the average gradients on labeled ID data, uncertainty score.
$e$ randomly generated unit vector.
$y$ target unit vector $e$ for ID node representations.
$\widehat{y}_{\mathbf{x}}$ predicted vector for input $\mathbf{x}$.
$h_{\theta_{t}}$ predictor on labeled in-distribution
$\boldsymbol{X}_{\text{wild}}^{\text{in}},\boldsymbol{X}_{\text{wild}}^{\text{out}}$ inliers and outliers in the wild dataset.
$\boldsymbol{X}^{\text{in}},\boldsymbol{X}_{\text{wild}}$ labeled ID data and unlabeled wild data.
$n,m$ size of $\boldsymbol{X}^{\text{in}}$, size of $\boldsymbol{X}_{\text{wild}}$
$T$ the filtering threshold
$\boldsymbol{X}_{T}$ wild data whose uncertainty score higher than threshold $T$
Distances
$r_{1}$ the radius of the hypothesis spaces $\mathcal{W}$
$\parallel\cdot\parallel_{2}$ $\ell_{2}$ norm
Loss, Risk and Predictor
$\ell(\cdot,\cdot)$ ID loss function
$R_{\boldsymbol{X}}(h_{\theta_{t}})$ the empirical risk w.r.t. predictor $h_{\theta_{t}}$ over data $\boldsymbol{X}$
$R_{\mathbb{P}_{\boldsymbol{XY}}}(h_{\theta_{t}})$ the risk w.r.t. predictor $h_{\theta_{t}}$ over distribution $\mathbb{P}_{\boldsymbol{XY}}$.
$ERR_{\text{out}}$ the error rate of regarding OOD as ID.
Table 6: Table of Notations and Descriptions
A.2 Definitions
Definition A.1.
($\beta$smooth).We say a loss function $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ (defined over $\boldsymbol{X}\times\boldsymbol{Y})$ is $\beta$smooth, if forany $\mathbf{x}\in{\boldsymbol{X}}$ and $y\in\boldsymbol{Y}$
$\begin{Vmatrix}\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)-\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\end{Vmatrix}_{2}\leq\beta\|\mathbf{w}-\mathbf{w}^{\prime}\|_{2}$
Definition A.2.
(Gradient-based Distribution Discrepancy). Given distributions $\mathbb{P}$ and $\mathbb{Q}$ defined over $X$ , the Gradient-based Distribution Discrepancy w.r.t. predictor $\mathbf{f}_{\mathrm{w}}$ and loss $t$ is
$d_{\mathbf{w}}^{\ell}(\mathbb{P},\mathbb{Q})=\left\|\nabla R_{\mathbb{P}}(h_{\theta_{t}},\widehat{h}_{\theta})-\nabla R_{\mathbb{Q}}(h_{\theta_{t}},\widehat{h}_{\theta})\right\|_{2},$
where $\widehat{h}_{\theta}$ is a classifier which returns the closest one-hot vector of $h_{\mathrm{w}}$: $R_{\mathbb{P}}(h_{\theta_{t}},\widehat{h}_{\theta})=\mathbb{E}_{\mathbf{x}\sim\mathbb{P}}\ell(h_{\theta_{t}},\widehat{h}_{\theta})$ and $R_{\mathbb{Q}}(h_{\theta_{t}},\widehat{h}_{\theta})=\mathbb{E_{\mathbf{x}\sim\mathbb{Q}}}\ell(h_{\theta_{t}},\widehat{h}_{\theta})$
Definition A.3.
$(\gamma,\xi)$discrepancy). We say a wild distribution $\mathbb{P}_{wild}$ has $(\gamma,\xi)$discrepancy w.r.t. an ID joint distribution $\mathbb{P}_{in}$ n, if $\gamma>\min_{\mathbf{w}\in\mathcal{W}}$ $R_{\mathbb{P}_{XY}}(h_{\theta})$ and for any parameter $\mathbf{w}\in\mathcal{W}$ satisfying that $R_{\mathbb{P},\boldsymbol{X}\boldsymbol{Y}}(h_{\theta_{t}})\leq\gamma$ should meet the following condition
$d_{\mathbf{w}}^{\ell}(\mathbb{P}_{in},\mathbb{P}_{wild})>\xi,$
where $R_{\mathrm{P}_{XY}}(h_{\theta})=\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{P}_{XY}}\ell(h_{\theta}(\mathbf{x}),y)$
A.3 Assumptions
Assumption 1.
  • The parameter space $\mathcal{W}\subset B(\mathbf{w}_{0},r_{1})\subset\mathbb{R}^{d}\left(\ell_{2}\right.$ ball of radius $r_{1}$ around $W_{0}$);
  • $\ell(h_{\theta_{t}}(\mathbf{x}),y)\geq 0$ and $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ is $\beta_{1}$smooth;
  • $\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\|\nabla\ell(h_{\theta_{0}}(\mathbf{x}),y)\|_{2}=b_{1};$
  • $\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\ell(h_{\theta_{0}}(\mathbf{x}),y)=B_{1}.$
Assumption 2. $\ell(\mathbf{f}(\mathbf{x}),\widehat{y}_{\mathbf{x}})\leq\min_{y\in\boldsymbol{Y}}\ell(\mathbf{f}(\mathbf{x}),y)$ , where $\widehat{y}_{\mathbf{x}}$ returns the closest vector of the predictor $\mathbf{f}$’s output on $\mathbf{x}$
A.4 Constants in Theory
Constants Description
$M=\beta_{1}r_{1}^{2}+b_{1}r_{1}+B_{1}$ the upper bound of loss $\ell(h_{\theta_{t}}(\mathbf{x}),y)$.
$M^{\prime}=2(\beta_{1}r_{1}+b_{1})^{2}$ the upper bound of gradient-based filtering score (Du et al., 2024a)
$\widehat{M}_{t}=(\sqrt{M^{\prime}/2}+1)/(2t)$ the upper bound of our resonance-based filtering score $\tau$ at the $t$-th epoch
$\tilde{M}=\beta_{1}M$ a constant for simplified representation
$d$ the dimensions of parameter spaces $\mathcal{W}$
$R_{\text{in}}^{*}$ the optimal ID risk, i.e., $R_{in}^{*}=\mathrm{min}_{\mathbf{w}\in\mathcal{W}}\mathbb{E_{\mathbf{x}\sim\mathbb{P}_{\mathrm{in}}}}\mathcal{L}_{1}(\mathbf{x},e)$
$\delta(T,t)$ the main error in 8
$\xi$ the discrepancy between $\mathbb{P}_{\text{in}}$ and $\mathbb{P}_{\text{wild}}$
$\pi$ the ratio of OOD distribution in $\mathbb{P}_{\text{wild}}$
$\alpha$ learning rate
Table 7: Constants in theory.
Appendix B Main Theorems
Theorem B.1.
If Assumptions 1 and 2 hold, $\mathbb{P}_{wild}$ has $(\gamma,\xi)$discrepancy w.r.t. $\mathbb{P}_{xy}$ ,and there exists $\eta\in(0,1)$ s.t. $\Delta=(1-\eta)^{2}\xi^{2}-8\beta_{1}R_{in}^{*}>0$, then for
$n=\Omega\big{(}\frac{\tilde{M}+M(r_{1}+1)d}{\eta^{2}\Delta}+\frac{M^{2}d}{(\gamma-R_{in}^{*})^{2}}\big{)},\quad m=\Omega\big{(}\frac{\tilde{M}+M(r_{1}+1)d}{\eta^{2}\xi^{2}}\big{)},$
with the probability at least 9/10 for any $0<T<\widehat{M}_{t}$ (here $\widehat{M}_{t}$ is the upper bound of filtering score $\tau_{i}$ at $t$-th epoch, i.e., $\tau_{i}\leq\widehat{M}_{t}$ )
$\text{ERR}_{\text{out}}^{t}\leq\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}+O(\sqrt{\frac{d}{\pi^{2}n}})+O(\sqrt{\frac{\mathrm{max}\{d,\Delta_{\xi}^{\eta^{2}}/\pi^{2}\}}{\pi^{2}(1-\pi)m}})$ (10)
where $\Delta_{\xi}^{\eta}=0.98\eta^{2}\xi^{2}-8\beta_{1}R_{in}^{*}$ and $R_{in}^{*}$ is the optimal ID risk, i.e., $R_{in}^{*}=\mathrm{min}_{\mathbf{w}\in\mathcal{W}}\mathbb{E_{\mathbf{x}\sim\mathbb{P}_{\mathrm{in}}}}\mathcal{L}_{1}(\mathbf{x},e)$. $d$ is the dimension of the space $\mathcal{W}$, $t$ denotes the $t$-th epoch, and $\pi$ is the OOD class-prior probability in the wild.
$M=\beta_{1}r_{1}^{2}+b_{1}r_{1}+B_{1},\quad\tilde{M}=M\beta_{1}$ (11)
Theorem B.2.
1) if $\Delta_{\xi}^{\eta}\geq(1-\epsilon)\pi$ for a small error $\epsilon\geq 0$, then the main error $\delta(T,t)$ satisfies that
$\delta(T,t)=\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\leq\frac{\epsilon}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}$ (12)
2) When learning rate $\alpha$ is small sufficiently, and if $\xi\geq 2.011\sqrt{8\beta_{1}R_{in}^{*}+1.011\sqrt{\pi}}$, then there exists $\eta\in(0,1)$ ensuring that $\Delta>0$ and $\Delta_{\xi}^{\eta}>\pi$ hold, which implies that the main error $\delta(T,t)=0$.
Appendix C Proofs of Main Theorems
C.1 Proof of Theorem 1
Step 1. With the probability at least $01-\frac{7}{3}\delta>$ 0
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim S_{\mathrm{wild}}^{\mathrm{in}}\tau_{i}}$ $\leq 8\beta_{1}R_{\mathrm{in}}^{*}$
$+4\beta_{1}\Big{[}C\sqrt{\frac{Mr_{1}(\beta_{1}r_{1}+b_{1})d}{n}}+C\sqrt{\frac{Mr_{1}(\beta_{1}r_{1}+b_{1})d}{(1-\pi)m-\sqrt{m\log(6/\delta)/2}}}$
$+3M\sqrt{\frac{2\log(6/\delta)}{n}}+M\sqrt{\frac{2\log(6/\delta)}{(1-\pi)m-\sqrt{m\log(6/\delta)/2}}}\Big{]},$
This can be proven by Lemma 7 in (Du et al., 2024a) and following inequality
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\mathcal{S}_{wild}^{in}}\tau_{i}\geq\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{wild}^{m}}\left\|\nabla\ell(h_{\theta_{\boldsymbol{X}^{m}}}(\tilde{\mathbf{x}}_{i}),\widehat{{h}}_{\theta_{\boldsymbol{X}^{m}}}(\tilde{\mathbf{x}}_{i}))-\mathbb{E}_{(\mathbf{x}_{j},y_{j})\sim\boldsymbol{X}^{m}}\nabla\ell(h_{\theta_{\boldsymbol{X}^{m}}}(\mathbf{x}_{j}),y_{j})\right\|_{2}^{2},$
Step 2.It is easy to check that
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}}\tau_{i}=\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{|\boldsymbol{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}}\tau_{i}+\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}{|\boldsymbol{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}}\tau_{i}.$
Step 3.Let
$\begin{gathered}\epsilon(n,m)=4\beta_{1}[C\sqrt{\frac{Mr_{1}(\beta_{1}r_{1}+b_{1})d}{n}}+C\sqrt{\frac{Mr_{1}(\beta_{1}r_{1}+b_{1})d}{(1-\pi)m-\sqrt{m\log(6/\delta)/2}}}\\ +3M\sqrt{\frac{2\log(6/\delta)}{n}}+M\sqrt{\frac{2\log(6/\delta)}{(1-\pi)m-\sqrt{m\log(6/\delta)/2}}}].\end{gathered}$
Under the condition in Theorem 5 in (Du et al., 2024a), with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}\tau_{i}}$ $\leq\frac{m}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big{[}\frac{98\eta^{2}\xi^{2}}{100}-\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}8\beta_{1}R_{\mathrm{in}}^{*}-\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}\epsilon(n,m)\big{]}$
$\leq\frac{m}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big{[}\frac{98\eta^{2}\xi^{2}}{100}-8\beta_{1}R_{\mathrm{in}}^{*}-\epsilon(n,m)\big{]}$
$\leq[\frac{1}{\pi}-\frac{\sqrt{\log 6/\delta}}{\pi^{2}\sqrt{2m}+\pi\sqrt{\log(6/\delta)}}\Big{]}\Big{[}\frac{98\eta^{2}\xi^{2}}{100}-8\beta_{1}R_{\mathrm{in}}^{*}-\epsilon(n,m)\Big{]}.$
In this proof, we set
$\Delta(n,m)=\big{[}\frac{1}{\pi}-\frac{\sqrt{\log 6/\delta}}{\pi^{2}\sqrt{2m}+\pi\sqrt{\log(6/\delta)}}\big{]}\big{[}\frac{98\eta^{2}\xi^{2}}{100}-8\beta_{1}R_{\mathrm{in}}^{*}-\epsilon(n,m)\big{]}.$
Note that $\Delta_{\xi}^{\eta}=0.98\eta^{2}\xi^{2}-8\beta_{1}R_{\mathrm{in}}^{*}$ , then
$\Delta(n,m)=\frac{1}{\pi}\Delta_{\xi}^{\eta}-\frac{1}{\pi}\epsilon(n,m)-\Delta_{\xi}^{\eta}\epsilon(m)+\epsilon(n)\epsilon(n,m),$
where $\epsilon(m)=\sqrt{\log 6/\delta}/(\pi^{2}\sqrt{2m}+\pi\sqrt{\log(6/\delta)}).$
Step 4. Under the conditions in Theorem 5 in (Du et al., 2024a) and Proposition D.4, with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{1-\min\{1,\Delta(n,m)\}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}},$ (13)
We prove this step: let $Z$ be the uniform random variable with $S_{\mathrm{wild}}^{\mathrm{out}}$ as its support and $Z(i)=$ $\tau_{i}/(\frac{\sqrt{2}}{2t\alpha-1})^{2}$ , then by the Markov inequality, we have
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}<T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}=P(Z(I)<T/(\frac{\sqrt{2}}{2t\alpha-1})^{2})\geq\frac{\Delta(n,m)-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}.$ (14)
Step 5. If $\pi\leq\Delta_{\xi}^{\eta}/(1-\epsilon/M^{\prime})$ , then with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{\epsilon+(\frac{\sqrt{2}}{2t\alpha-1})^{2}\epsilon^{\prime}(n,m)}{(\frac{\sqrt{2}}{2t\alpha-1})^{2}-T},$ (15)
where $\epsilon^{\prime}(n,m)=\epsilon(n,m)/\pi+\Delta_{\xi}^{\eta}\epsilon(m)-\epsilon(n)\epsilon(n,m).$
Step 6. If we set $\delta=3/100$ , then it is easy to see that
$\epsilon(m)\leq O({\frac{1}{\pi^{2}\sqrt{m}}}),$
$\epsilon(n,m)\leq O(\beta_{1}M\sqrt{\frac{d}{n}})+O(\beta_{1}M\sqrt{\frac{d}{(1-\pi)m}}),$
$\epsilon^{\prime}(n,m)\leq O(\frac{\beta_{1}M}{\pi}\sqrt{\frac{d}{n}})+O\Big{(}(\beta_{1}M\sqrt{d}+\sqrt{1-\pi}\Delta_{\xi}^{\eta}/\pi)\sqrt{\frac{1}{\pi^{2}(1-\pi)m}}\Big{)}.$
Step 7. By results in Steps 4, 5 and 6, We complete this proof
C.2 Proof of Theorem 2
The first result is trivial. Hence,we omit it.We mainly focus on the second result in this theorem In this proof, then we set
$\eta=\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}+0.99\pi}/(\sqrt{0.98}\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}}+\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}+\pi})$
Note that it is easy to check that
$\xi\geq 2.011\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}}+1.011\sqrt{\pi}\geq\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}}+1.011\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}+\pi}.$
Therefore,
$\eta\xi\geq\frac{1}{\sqrt{0.98}}\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}+0.99\pi}>\sqrt{8\beta_{1}R_{\mathrm{in}}^{*}+\pi},$
which implies that $\Delta_{\xi}^{\eta}>\pi$ Note that
$(1-\eta)\xi\geq\frac{1}{\sqrt{0.98}}\big{(}\sqrt{0.98}\sqrt{8\beta_{1}R_{\mathrm{m}}^{*}}+\sqrt{8\beta_{1}R_{\mathrm{m}}^{*}+\pi}-\sqrt{8\beta_{1}R_{\mathrm{m}}^{*}+0.99\pi}\big{)}>\sqrt{8\beta_{1}R_{\mathrm{m}}^{*}},$
which implies that $\Delta>0$ We have completed this proof
Appendix D Necessary Propositions.
D.1 Boundedness
Proposition D.1.
If Assumption 1 holds,
$\begin{gathered}\operatorname*{sup}_{\mathbf{w}\in\mathcal{W}}\operatorname*{sup}_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\|\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\|_{2}\leq\beta_{1}r_{1}+b_{1}={\sqrt{M^{\prime}/2}},\\ \sup_{\mathbf{w}\in\mathcal{W}}\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\ell(h_{\theta_{t}}(\mathbf{x}),y)\leq\beta_{1}r_{1}^{2}+b_{1}r_{1}+B_{1}=M,\\ \end{gathered}$
Proof. One can prove this by Mean Value Theorem of Integrals easily.
Proposition D.2.
If Assumption 1 holds, for any $\mathbf{w}\in\mathcal{W}$,
$\left\|\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\right\|_{2}^{2}\leq 2\beta_{1}\ell(h_{\theta_{t}}(\mathbf{x}),y).$
Proof. The details of the self-bounding property can be found in Appendix B of Lei Ying
Proposition D.3.
If Assumption 1 holds, for any labeled data $\boldsymbol{X}$ and distribution $\mathbb{P}$.
$\left\|\nabla R_{\boldsymbol{X}}(h_{\theta_{t}})\right\|_{2}^{2}\leq 2\beta_{1}R_{\boldsymbol{X}}(h_{\theta_{t}}),\quad\forall\mathbf{w}\in\mathcal{W},$ (16)
$\left\|\nabla R_{\mathbb{P}}(h_{\theta_{t}})\right\|_{2}^{2}\leq 2\beta_{1}R_{\mathbb{P}}(h_{\theta_{t}}),\quad\forall\mathbf{w}\in\mathcal{W}.$ (17)
Proof. Jensen’s inequality implies that $R_{S}(h_{\theta_{t}})$ and $R_{\mathbb{P}}(\mathbf{f}_{\mathrm{w}})$ are $\beta_{1}$smooth.Then Proposition 2 implies the results.
Proposition D.4.
If Assumption 1 holds, for any $\mathbf{w}_{t}\in\mathcal{W}$,
$\parallel\Delta h_{\theta_{t}}({\mathbf{x}})\parallel_{2}\leq(\sqrt{M^{\prime}/2}+1)/(2t)=\widehat{M}_{t}$
Proof. It is trivial that
$\parallel\mathbf{x}^{\top}\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\parallel\leq\parallel\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\parallel\leq\beta_{1}r_{1}+b_{1}={\sqrt{M^{\prime}/2}}$
Then
$\parallel\mathbf{x}^{\top}\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\parallel=\parallel 2(\mathbf{x}\mathbf{W}^{\top}-y)\parallel\geq 2\parallel\sum_{t}\Delta h_{\theta_{t}}(\mathbf{x})-y\parallel\geq 2\parallel t\Delta h_{\theta_{t}}(\mathbf{x})-y\parallel\geq 2t\parallel\Delta h_{\theta_{t}}(\mathbf{x})\parallel-1$
It is straightforward to verify that:
$\|\Delta h_{\theta_{t}}(\mathbf{x})\|_{2}\leq\frac{\sqrt{M^{\prime}/2}+1}{2t}\leq\alpha\sqrt{M^{\prime}/2}=\widehat{M}_{t}.$
Here, $\alpha$ is the learning rate. From the inequality above, we establish a relationship between $\sqrt{M^{\prime}/2}$, $\alpha$, and $t$ as follows:
$M^{\prime}\geq(\frac{\sqrt{2}}{2t\alpha-1})^{2}.$
Appendix E A Straightforward Explanation of Feature Resonance
To verify the phenomenon of Feature Resonance, we calculate the change $\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})$ in the representation $h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})$ of an unlabeled node $i$ from the $t$-th ($t\geq 0$) epoch to the $(t+1)$-th epoch, defined as follows:
$\begin{split}&\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\\ &=h_{\theta_{t+1}}(\tilde{\mathbf{x}}_{i})-h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\\ &=-\alpha\ {\tilde{\mathbf{x}}_{i}\ \nabla_{\theta_{t}}\ell({\boldsymbol{X}}_{\text{known}})}\\ &=2\alpha\mathbb{E}(\underbrace{\tilde{\mathbf{x}}_{i}{\boldsymbol{X}}_{\text{known}}^{\top}}_{\text{Term 1}}(\underbrace{({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top}_{t})-\mathbf{1}^{\top}e)}_{\text{Term 2}})\end{split}$ (18)
where $\alpha$ is the learning rate. Term 1 in the Equation 18 illustrates that when the features of $\tilde{\mathbf{x}}_{i}$ are consistent with the overall features of the labeled ID nodes ${\boldsymbol{X}}_{\text{known}}$, the representation of $\tilde{\mathbf{x}}_{i}$ undergoes a more significant change. Meanwhile, since term 2 in the Equation 18 and $\tilde{\mathbf{x}}_{i}$ are independent, the choice of the target vector can be arbitrary. It is highly suitable for category-free OOD detection scenarios, requiring no multi-category labels as ground truth.
Appendix F Experiment Details
We supplement experiment details for reproducibility. Our implementation is based on Ubuntu 20.04, Cuda 12.1, Pytorch 2.1.2, and Pytorch Geometric 2.6.1. All the experiments run with an NVIDIA 3090 with 24GB memory.
Table 8: Hyper-parameters for training.
Dataset Squirrel WikiCS YelpChi Amazon Reddit
Learning rate ($\alpha$) 0.005 0.01 0.005 0.005 0.01
$h_{\theta}$ layers 1 1 1 1 1
$g_{\theta}(\cdot)$ layers 2 2 2 2 2
Hidden states 16 16 16 16 16
Dropout rate 0.1 0.1 0.1 0.1 0.1
$\mathrm{n}$ 2 1 2 2 1
$\lambda$ 0.5 0.5 0.5 0.5 0.5
F.1 Hyperparameter
As shown in Table 8.
F.2 Metric
Following prior research on OOD node detection, we evaluate the detection performance using three widely recognized, threshold-independent metrics: AUROC ($\uparrow$), AUPR ($\uparrow$) and FPR95($\downarrow$). (1) AUROC measures the area under the receiver operating characteristic curve, capturing the trade-off between the true positive rate and the false positive rate across different threshold values. (2) AUPR calculates the area under the precision-recall curve, representing the balance between the precision rate and recall rate for OOD nodes across varying thresholds. (3) FPR95 is defined as the probability that an OOD sample is misclassified as an ID node when the true positive rate is set at 95%.
F.3 Dataset Description
To thoroughly evaluate the effectiveness of RSL, we perform experiments on five diverse and real-world OOD node detection datasets:
  • Squirrel (Rozemberczki et al., 2021): A Wikipedia network where nodes correspond to English Wikipedia articles, and edges represent mutual hyperlinks. Nodes are categorized into five classes following Geom-GCN (Pei et al., 2020) annotations, with the network exhibiting a high level of heterophily.
  • WikiCS (Mernyei & Cangea, 2020): This dataset consists of nodes representing articles in the Computer Science domain. Edges are based on hyperlinks, and nodes are classified into 10 categories, each corresponding to a unique sub-field of Computer Science.
  • YelpChi (Rayana & Akoglu, 2015): Derived from Yelp, this dataset includes hotel and restaurant reviews. Legitimate reviews are labeled as ID nodes, while spam reviews are considered OOD nodes.
  • Amazon (McAuley & Leskovec, 2013): Contains reviews from the Musical Instrument category on Amazon.com. ID nodes represent benign users, while OOD nodes correspond to fraudulent users.
  • Reddit (Kumar et al., 2019): A dataset comprising user posts collected from various subreddits over a month. Normal users are treated as ID nodes, while banned users are labeled as OOD nodes.
We follow the same data preprocessing steps as EnergyDef (Gong & Sun, 2024). Both Squirrel and WikiCS datasets are loaded using the DGL (Wang et al., 2019) package. For Squirrel, class {1} is selected as the OOD class, while {0, 2, 3, 4} are designated as ID classes. In the case of WikiCS, {4, 5} are chosen as OOD classes, with the remaining eight classes treated as ID. The YelpChi and Amazon datasets are processed based on the methodology described in (Dou et al., 2020), and the Reddit dataset is prepared using the PyGod (Liu et al., 2022) package.
F.4 Baseline Description
  • LOF-KNN (Breunig et al., 2000) calculates the OOD scores of node attributes by assessing the deviation in local density relative to the k-nearest node attributes.
  • MLPAE uses an MLP-based autoencoder, where the reconstruction error of node attributes is used as the OOD score. It is trained by minimizing the reconstruction error on ID training nodes.
  • GCNAE (Kipf & Welling, 2016b) swaps the MLP backbone for a GCN in the autoencoder. The OOD score is determined in the same way as MLPAE, following the same training process.
  • GAAN (Chen et al., 2020) is a generative adversarial network for attributes that evaluates sample reconstruction error and the confidence of recognizing real samples to predict OOD nodes.
  • DOMINANT (Ding et al., 2019) combines a structure reconstruction decoder and an attribute reconstruction decoder. The total reconstruction error for each node consists of the errors from both decoders.
  • ANOMALOUS (Peng et al., 2018) is an anomaly detection method that utilizes CUR decomposition and residual analysis for identifying OOD nodes.
  • SL-GAD (Zheng et al., 2021) derives OOD scores for nodes by considering two aspects: reconstruction error and contrastive scores.
  • GOAD (Bergman & Hoshen, 2020) enhances training data by transforming it into independent spaces and trains a classifier to align the augmented data with the corresponding transformations. OOD scores are then calculated based on the distances between OOD inputs and the centers of the transformation spaces. For graph-structured data, we use the same GNN backbone as EnergyDef-h.
  • NeuTral AD (Qiu et al., 2021) uses learnable transformations to embed data into a semantic space. The OOD score is determined by a contrastive loss applied to the transformed data.
  • GKDE (Zhao et al., 2020) predicts Dirichlet distributions for nodes and derives uncertainty as OOD scores by aggregating information from multiple sources.
  • OODGAT (Song & Wang, 2022) is an entropy-based OOD detector that assumes node category labels are available. It uses a Graph Attention Network as the backbone and determines OOD nodes based on category distribution outcomes.
  • GNNSafe (Wu et al., 2023) calculates OOD scores by applying the LogSumExp function over the output logits of a GNN classifier, which is trained with multi-category labels. The rationale for the OOD score is the similarity between the Softmax function and the Boltzmann distribution.
  • SSD (Sehwag et al., 2021) is an outlier detector that leverages self-supervised representation learning and Mahalanobis distance-based detection on unlabeled ID data. We use twice dropout to generate positive pairs for contrastive learning like SimCSE (Gao et al., 2021).
  • EnergyDef (Gong & Sun, 2024) uses Langevin dynamics to generate synthetic OOD nodes for training the OOD node classifier.
F.5 Implementation Details
We adopt the same dataset settings as EnergyDef (Gong & Sun, 2024). It is worth noting that, under this dataset setup, the features of unknown nodes are accessible. Therefore, using the features of unknown nodes during the training phase to filter reliable OOD nodes is a legitimate strategy. Specifically, for the Squirrel and WikiCS datasets, we randomly select one and two classes as OOD classes, respectively. In the case of fraud detection datasets, we categorize a large number of legitimate entities as ID nodes and fraudsters as OOD nodes. We allocate 40% of the ID class nodes for training, with the remaining nodes split into a 1:2 ratio for validation and testing, ensuring stratified random sampling based on ID/OOD labels.
We report the average value of five independent runs for each dataset. The hyper-parameters are shown in Table 8. The anomaly detection baselines are trained entirely based on graph structures and node attributes without requiring ID annotations. We adapt these models to the specifications of our OOD node detection tasks by minimizing the corresponding loss items solely on the ID nodes, where applicable.
Appendix G More Experiments
(a) Reddit
(b) Squirrel
(c) YelpChi
(d) WikiCS
Figure 6: The performance of using resonance-based score $\tau$ to detect OOD nodes varies with training progress. The higher the AUROC, the better, and the lower the FPR95, the better.
G.1 The Feature Resonance Phenomenon Induced by Different Target Vectors
We explore the phenomenon of feature resonance using different target vectors. Experiments are conducted on two datasets with real $N$-category labels, Squirrel and WikiCS ($N$ represents the number of categories). First, based on the neural collapse theory (Papyan et al., 2020; Zhou et al., 2022), we preset $N$ target vectors, each representing a category. These $N$ target vectors form an equiangular tight frame, maximizing the separation between them. The definition of the simplex equiangular tight frame is introduced as follows:
Definition G.1.
Simplex ETF. (Xiao et al., 2024) A simplex equiangular tight frame (ETF) refers to a collection of K equal-length and maximally-equiangular P-dimensional embedding vectors $\mathbf{E}=[e_{1},\cdots,e_{K}]\in\mathbb{R}^{P\times K}$ which satisfies:
$\mathbf{E}=\sqrt{\frac{K}{K-1}}\mathbf{U}\big{(}\mathbf{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\big{)}$ (19)
where $\mathbf{I}_{K}$ is the identity matrix,$\mathbf{1}_{K}$ is an all-ones vector, and $\mathbf{U}\in\mathbb{R}^{P\times K}(P\geq K)$ allows a rotation.
All vectors in a simplex ETF $\mathbf{E}$ have an equal $\mathit{l}_{2}$ norm and the same pair-wise maximal equiangular angle $-\frac{1}{K-1}$,
$e_{k_{1}}^{\top}e_{k_{2}}=\frac{K}{K-1}\delta_{k_{1},k_{2}}-\frac{1}{K-1},\forall k_{1},k_{2}\in[1,K]$ (20)
where $\delta_{k_{1},k_{2}}=1$ when $k_{1}=k_{2}$ and 00 otherwise.
We use MSE loss to pull the representations of known ID nodes toward their corresponding target vectors based on their labels, as follows:
$\ell(h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{E}_{\text{known}}-({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (21)
where $\mathbf{E}_{\text{known}}$ denotes the target vector matrix corresponding to the known ID nodes.
The trajectory trends and lengths of unknown ID nodes differ significantly from those of OOD nodes, with the former showing more distinct trends and longer trajectories. We refer to this as the feature resonance phenomenon and leverage it to filter OOD nodes. As shown in Table 4, under the “True multi-label” row, the experimental results demonstrate that this method is effective and performs well. Interestingly, even with random labels for known ID nodes or aligning all known ID representations to a fixed target vector, unknown ID nodes consistently exhibit longer trajectories than unknown OOD nodes, as shown in Table 4.
The experiments above indicate that the feature resonance phenomenon is label-independent and results from the intrinsic relationships between ID node representations. Therefore, this is highly suitable for category-free OOD detection scenarios without multi-category labels.
(a) WikiCS
(b) YelpChi
Figure 7: The impact of different sliding window widths on the performance of detecting OOD nodes. When the width is 1, it corresponds to the resonance-based score $\tau$.
(a) Pre-training (EnergyDef)
(b) Pre-training (Ours)
(c) Post-training (EnergyDef)
(d) Post-training (Ours)
Figure 8: T-SNE visualization of node embeddings on the dataset WikiCS. (a) Synthetic nodes (red) generated by EnergyDef fail to accurately represent the actual features of OOD nodes (blue), whereas ours can, as shown in (b). (c) Representations of ID (green) and OOD (blue) nodes trained with synthetic nodes generated by EnergyDef are poorly separated, whereas ours can, as shown in (d).
G.2 Variation of Microscopic Feature Resonance During Training
We also observe the variation of the microscopic feature resonance phenomenon during the training process on other datasets, as shown in Figure 6. We find that the changes on Reddit, YelpChi, and WikiCS are generally consistent with Amazon, with the most significant feature resonance occurring in the middle of the training process. However, for Squirrel, the feature resonance phenomenon reaches its most pronounced level early in the training. We believe this is due to the relatively rich features in Squirrel, which allow the model to quickly identify the optimal optimization path for ID samples in the early stage of training.
G.3 The Impact of Different Sliding Window Widths on Performance
We investigate the impact of different sliding window widths on the effectiveness of detecting OOD nodes. The experimental results in Figure 7 show that as the sliding window width increases, the detection performance for OOD nodes gradually decreases. This suggests that a more fine-grained observation leads to better detection performance.
G.4 Node Representation Visualization
EnergyDef generates auxiliary synthetic OOD nodes via SGLD to train an OOD classifier for category-free OOD node detection. However, we find that the synthetic OOD nodes from EnergyDef do not accurately capture the features of actual OOD nodes. As shown in Figure 8(a), most synthetic OOD nodes are separated from actual OOD nodes and even overlap with ID nodes, limiting the classifier’s performance. The severe overlap between ID and OOD node representations after training by EnergyDef (Figure 8(c)) further highlights this issue. In contrast, we use feature resonance to identify reliable OOD nodes and synthesize new ones based on these. As seen in Figure 8(b), our synthetic OOD nodes align more closely with the actual OOD nodes. Training with these nodes results in better separation between ID and OOD node representations, as shown in Figure 8(d).
Appendix H Discussion
H.1 Differences from Gradient-Based Methods
It is important to note that our method RSL differs significantly from previous gradient-based methods:
1) Originating from the Commonality of Representations. Our method is based on the conjecture that there are inherent commonalities between the representations of the ID sample, which are independent of gradients.
2) No Pre-trained Multi-category Classifier Required. Gradient-based methods like GradNorm (Huang et al., 2021) compute the KL divergence between an unknown sample’s softmax output from a multi-category classifier and a uniform distribution, using the gradient norm to distinguish OOD samples. OOD samples, with uniform softmax outputs, yield more minor gradient norms, whereas sharper outputs for ID samples produce more significant norms. Similarly, SAL (Du et al., 2024a) uses pseudo-labels from a multi-category classifier for unknown samples, continuing training to compute gradients, and identifies OOD samples via the gradient’s principal component projection. These methods require a pre-trained multi-category classifier, making them unsuitable for category-free scenarios without labels, whereas our RSL method avoids this limitation.
3) No Need to Compute Gradients for Unknown Samples. As shown in Equation 18, we only need the representations of unknown samples to compute our resonance-based score. This significantly enhances the flexibility of our method, as we can detect OOD samples during any optimization of known ID representations without the need to wait until after the optimization is complete.
Appendix I Algorithm Pseudo-code
Algorithm 1 Resonance-based Separate and Learn (RSL) Framework for Category-Free OOD Detection
1:  Input: Known ID nodes $\mathcal{V}_{\mathrm{known}}$, Wild nodes $\mathcal{V}_{\mathrm{wild}}$, Target vector $e$ with random initial, Validation set $\mathcal{V}_{\mathrm{val}}$
2:  Output: OOD classifier $E_{\theta}$
3:  Phase 1: Feature Resonance Phenomenon
4:  Initialize model $h_{\theta}$ with random parameters $\theta$
5:  for $t=1$ to $\mathbb{T}$ (training epochs) do
6:     Optimize $h_{\theta_{t}}(\cdot)$ to align $\mathcal{V}_{\mathrm{known}}$ with target $e$:
$\ell(h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{1}^{\top}e-({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$
7:     Calculate the representation change of $\tilde{v}_{i}\in\mathcal{V}_{\mathrm{wild}}$ : $\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})=h_{\theta_{t+1}}(\tilde{\mathbf{x}}_{i})-h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})$
8:     Compute resonance-based score $\tau_{i}=\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{2}$
9:  end for
10:  Identify the period of feature resonance using the validation set, selecting $t$ where $\tau$ best separates ID and OOD nodes.
11:  Phase 2: Candidate OOD Node Selection
12:  Define candidate OOD set:
$\mathcal{V}_{\mathrm{cand}}=\{\tilde{v}_{i}\in\mathcal{V}_{\mathrm{wild}}:\tau_{i}\leq T\}$
13:  Phase 3: Synthetic OOD Node Generation
14:  for each $\hat{v}_{j}\in\mathcal{V}_{\mathrm{syn}}$ (synthetic OOD nodes) do
15:     Generate $\hat{\mathbf{x}}_{j}^{(t+1)}$ with random initial using:
$\hat{\mathbf{x}}_{j}^{(t+1)}=\lambda\big{(}\hat{\mathbf{x}}_{j}^{(t)}-\frac{\alpha}{2}\nabla_{\hat{\mathbf{x}}_{j}^{(t)}}E_{\theta}(\hat{v}_{j}^{(t)})+\epsilon\big{)}+(1-\lambda)\mathbb{E}_{\mathbf{x}\sim\boldsymbol{X}_{\mathrm{cand}}}(\mathbf{x}-\hat{\mathbf{x}}_{j}^{(t)}),,\epsilon\sim\mathcal{N}(0,\zeta)$
16:  end for
17:  Phase 4: OOD Classifier Training
18:  Define training set $\mathcal{V}_{\mathrm{train}}=\mathcal{V}_{\text{known}}\cup\mathcal{V}_{\mathrm{cand}}\cup\mathcal{V}_{\mathrm{syn}}$
19:  Assign labels $\boldsymbol{Y}_{\mathrm{train}}$ for ID nodes ($1$) and OOD nodes (00)
20:  Train $E_{\theta}$ using binary cross-entropy loss:
$\ell_{\text{cls}}=\mathbb{E}_{v\sim\mathcal{V}_{\mathrm{train}}}\big{(}\mathrm{y}_{v}\mathrm{log}(\sigma(E_{\theta}(v)))+(1-\mathrm{y}_{v})\mathrm{log}(1-\sigma(E_{\theta}(v)))\big{)}$
21:  Return: Trained OOD classifier $E_{\theta}$